A rectangle is 15 inches longer than its wide. Its perimeter is 150 inches. What is the rectangles area
Answer by JulietG(1812)

P = 150
L = W + 15
P = 2L + 2W
150 = 2L + 2W
150 = 2(W+15) + 2W
150 = 2W + 30 + 2W
150 = 4W + 30
120 = 4W
30 = W
Therefore, L = 45
A = L * W
A = 30 * 45
Area = 1350